Two protons are separated by a distance of 2.0×10-15. What is the magnitude of the electric force between them
1
Expert's answer
2020-05-04T12:46:38-0400
"F=k\frac{e^2}{r^2}=(9\cdot 10^{9})\frac{(1.6\cdot 10^{-19})^2}{(2\cdot 10^{-15})^2}=58\ N"